# SACNet: A Spatially Adaptive Convolution Network for 2D Multi-organ Medical Segmentation thanks: *Corresponding author.This work is supported by National Natural Science Foundation of China [No. 62371156] and Natural Science Foundation of Guangdong Province [No. 2022A1515011629]

### v1

@@ V-A Performance on Synapse Dataset
On the Synapse dataset, as shown in Table I, our SACNet stands out with the highest scores for most organs, achieving 89.36% for the left kidney, 85.41% for the right kidney, 95.98% for the liver, and 87.64% for the stomach. For other organs, except the spleen, our SACNet also achieves the second-best scores. Additionally, our DSC reaches 84.92% and HD95 is 15.13, both being the best results compared to other benchmarks. Figure 3 provides visual examples of the segmentation outputs produced by our proposed model and other state-of-the-art methods. As shown, our predictions have clearer boundaries between different segmentation targets: the boundaries between the pancreas and spleen are clear and smooth. Our model is more accurate for small objects: Swin-UNet and AgileFormer miss the gallbladder in their segmentations, and TransUNet only partially segments it, but our model segments it completely. As shown in the last two rows of Table I, replacing CDLoss with CTLoss significantly improves the results for DSC and HD95, as well as for most organs, particularly the left and right kidneys. In most slice data, the left and right kidneys have complex boundaries and overlap with surrounding organs, making them relatively difficult to segment, as shown in the orange and cyan regions in Figure 3. This demonstrates that our CTLoss is highly effective in balancing easy and difficult classes.
V-B Performance on ACDC Dataset
On the ACDC dataset, our SACNet outperforms other models in terms of segmentation accuracy. As shown in Table II, our proposed SACNet achieves the best segmentation results compared to other benchmarks, with a Dice score of 92.53%, representing a significant improvement of 4.46% over Swin-UNet and a 2.82% increase over TransUNet. Furthermore, in Figure 3, we provide visual illustrations of the segmentation outputs produced by our proposed model and other state-of-the-art approaches. As demonstrated, our predictions have clearer boundaries between different segmentation targets and align closely with the provided ground truth masks, indicating the effectiveness of our approach. One notable advantage of SACNet is its capability to model adaptive receptive fields according to different segmentation targets. Additionally, the inclusion of our CTLoss results in a significant improvement in segmentation accuracy for difficult classes compared to easy classes, such as the right ventricle. In most slice data, the RV is smaller in volume and has a higher overlap with other organs, making it a difficult class, as shown in the red portion of Figure 3. This result demonstrates the effectiveness of our loss function in balancing categories.
V-C Ablation Study
Taking the Synapse dataset as an example, we ablate the key components in our SACNet to verify the effectiveness of these designs. As shown in Table III, we evaluate the effectiveness of our pre-trained DCNv3 blocks, FFN blocks, Layer Scale, and DropPath blocks respectively. When these key modules are removed, the model’s performance decreases to varying degrees. Without DCNv3, the model’s DSC drops from 84.92% to 76.99%, and HD95 increases from 15.13 to 31.05, with performance declining the most severely. This demonstrates that DCNv3 plays a vital role in our model. We also compare our model using two different loss functions: one combining cross-entropy and Dice loss, and the other using our proposed continuity dynamic adjustment loss function. The results show that using CTLoss for updating the model achieves better performance, and CTLoss is highly effective in balancing easy classes and difficult classes. Based on our comparisons, it is evident that the ARFM we designed, which integrates pre-trained DCNv3 blocks, FFN blocks, Layer Scale blocks, and DropPath blocks, can significantly enhance accuracy and establish SACNet as a new state-of-the-art model for the 2D multi-organ medical image segmentation task.
VI Conclusion
@@
In this study, we focus on the 2D multi-organ medical segmentation task and utilize the knowledge of deformable convolution v3 and multi-object segmentation to optimize our SACNet in three aspects: feature extraction, model architecture, and loss constraint, simultaneously enhancing perception for different segmentation targets. Firstly, we propose the Adaptive Receptive Field Module (ARFM), which combines (DCNv3) with a series of customized block-level and architecture-level designs similar to transformers. This module can capture the unique features of different organs by adaptively adjusting the receptive field according to various targets. Secondly, we utilize ARFM as building blocks to construct the encoder-decoder of SACNet and partially share parameters between the encoder and decoder, making the network wider rather than deeper, thereby achieving a more parameter-efficient and effective framework. Lastly, we propose a novel continuity dynamic adjustment loss function, based on t-vMF Dice loss and cross-entropy loss, to better balance easy and complex classes in segmentation. Our method is verified on 3D slice datasets from ACDC and Synapse, and the results show that our method provides better accuracy in multi-organ segmentation tasks compared to several existing methods.
However, there are limitations to our work compared to 3D segmentation methods. 3D segmentation can utilize the complete information from volumetric data, providing higher accuracy, but its computational and memory complexity is also significantly higher. In the future, we aspire to extend our method to 3D segmentation tasks to further enhance performance.

### v2

@@ IV-A Performance on Synapse Dataset
On the Synapse dataset, as shown in Table I, our SACNet stands out with the highest scores for most organs, achieving 89.36% for the left kidney, 85.41% for the right kidney, 95.98% for the liver, and 87.64% for the stomach. For other organs, except the spleen, our SACNet also achieves the second-best scores. Additionally, our DSC reaches 84.92% and HD95 is 15.13, both being the best results compared to other benchmarks. Figure 2 provides visual examples of the segmentation outputs produced by our proposed model and other state-of-the-art methods. As shown, our predictions have clearer boundaries between different segmentation targets: the boundaries between the pancreas and spleen are clear and smooth. Our model is more accurate for small objects: Swin-UNet and AgileFormer miss the gallbladder in their segmentations, and TransUNet only partially segments it, but our model segments it completely. As shown in the last two rows of Table I, replacing CDLoss with CTLoss significantly improves the results for DSC and HD95, as well as for most organs, particularly the left and right kidneys.
IV-B Ablation Study
Using the Synapse dataset, we ablated SACNet’s key components to validate their effectiveness, as detailed in Table II. We evaluated the impact of removing pre-trained DCNv3, FFN, Layer Scale, and DropPath blocks. Without DCNv3, DSC dropped from 84.92% to 76.99%, and HD95 rose from 15.13 to 31.05, indicating its crucial role. We also compared two loss functions: a combination of cross-entropy and Dice loss versus our proposed CTLoss, with the latter showing superior results. These comparisons confirm that our ARFM design, integrating key components, enhances accuracy, positioning SACNet as a state-of-the-art model for 2D multi-organ segmentation.
V Conclusion
@@
In this study, we focus on the 2D multi-organ medical segmentation task and utilize the knowledge of deformable convolution v3 and multi-object segmentation to optimize our SACNet in three aspects: feature extraction, model architecture, and loss constraint, simultaneously enhancing perception for different segmentation targets. Our method is verified on 3D slice datasets from Synapse, and the results show that our method provides better accuracy in multi-organ segmentation tasks compared to several existing methods.